OpenStudy (anonymous):

Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. I need help learning how to solve this problem I don't get it just step by step on how

OpenStudy (anonymous):

The first thing you have to do is pick numbers for the blanks.

OpenStudy (anonymous):

@peachpi do I just pick any numbers

OpenStudy (anonymous):

yes and for the 2nd blank decide whether you want it to increase or decrease

OpenStudy (anonymous):

Ok so if I say she practices piano for 10 on Monday and increases by 5 min everyday

OpenStudy (anonymous):

ok. so now you can write the formula for the sequence. \[a_n=a_1+d(n-1)\] \(a_1\)=10 because that was how much she practiced on the first day d = 5 because that's how much it increases each day. Plug them in to get the formula for you specific numbers

OpenStudy (anonymous):

did you get a formula?

OpenStudy (anonymous):

No how do I plug it in

OpenStudy (anonymous):

Oh never mind I see I'll write it down

OpenStudy (anonymous):

So an stays blank

OpenStudy (anonymous):

right. both an & n don't change for this step

OpenStudy (anonymous):

Ok I got an=15+5n

OpenStudy (anonymous):

So or idk how to write the a and the n the way u did my bad

OpenStudy (anonymous):

an = a1 + d(n-1) an = 10 + 5(n - 1) an = 10 + 5n - 5 an = 5n + 5 I think you may have added wrong somewhere

OpenStudy (anonymous):

Now for the last step you plug in 7 for n to find \(a_7\). \[a_7=5(7)+5\]

OpenStudy (anonymous):

I added the 5 to the 10

OpenStudy (anonymous):

ok. That was it. It was supposed to be -5 added to the 10

OpenStudy (anonymous):

So would it be 15 + 5n or no

OpenStudy (anonymous):

It's 5 + 5n

OpenStudy (anonymous):

Ok I got a7=40

OpenStudy (anonymous):

that's right :)

OpenStudy (anonymous):

Wait so that's it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Omg thank you so much could you help me with another if you can it's the same but with a geometric sequence

OpenStudy (anonymous):

yeah. just post another question and tag me